Unpolarised light is incident on two polarizing sheets placed over one another. The
intensity of light transmitted finally is one-third of the intensity of the incident light.
Calculate the angle between the transmission axes of the two polarizing sheets.
1
Expert's answer
2018-03-22T10:33:08-0400
The intensity of transmitted light for these two sheets at an angle θ is
I=1/2 I_0 cos^2⁡θ
Thus,
1/3 I_0=1/2 I_0 cos^2⁡θ
cos^2⁡θ=2/3
θ=cos^(-1)⁡√(2/3)=35.2°.